OpenStudy (anonymous):

Please Help... Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (johnnydicamillo):

\[y+4 =-3(x-3)\]

OpenStudy (anonymous):

I Don't Understand.

OpenStudy (anonymous):

Point slope form is this y – y1 = m(x – x1) where m is the slope

OpenStudy (johnnydicamillo):

Now just simplify the equation. Distribute the the "-3" and then subtract the 4 to both sides.

OpenStudy (anonymous):

y=-3x+5

OpenStudy (anonymous):

I subtract 4 from 5 and -3?

OpenStudy (anonymous):

no i gave you the answer, wait what lol

OpenStudy (anonymous):

She is trying to figure out how to do it

OpenStudy (anonymous):

Where did you get the 5 from?

OpenStudy (anonymous):

y-(-1)= -3 (x-3) y+4= -3x+9 *subtract the 4 and minus it to the 9* y=-3x+5

OpenStudy (anonymous):

i meant -4 not -1

OpenStudy (anonymous):

I'm extremely confused.. I've never done Algebra and my Geometry teacher never helped me so I have no idea how to do this.

OpenStudy (anonymous):

i showed you the work...

OpenStudy (johnnydicamillo):

Okay it is like this after you have the expression I gave you: \[y-y1 = slope(x - x1) where y1 and x1\] where x1 and y1 is your point. So when you plug it in you get the equation \[y+4 = -3(x-3)\] now you then distribute the "-3" so the equation looks like this: \[y+4 = -3x + 9\] now subtract the 4 so it looks like this \[y = -3x + 5\]

OpenStudy (johnnydicamillo):

sorry funky answer. disregard the very first line

OpenStudy (anonymous):

you just basically plug in your points, (x,y) and your slope which is m; to the equation y-y1=m(x-x1)

OpenStudy (johnnydicamillo):

yes

OpenStudy (anonymous):

Ahh I understand now, Thank You Guys!!

OpenStudy (anonymous):

welcome (: